For a normal probability distribution with a mean of 240 and a standard deviation of 40, calculate the following:
a. The probability that an observation will fall between 240 and 300.
b. The probability that an observation will fall between 195 and 240.
c. The probability that an observation will be less than 260.
d. The probability that an observation will fall between 180 and 220.

Answer by stanbon(75887)

For a normal probability distribution with a mean of 240 and a standard deviation of 40, calculate the following:
-------
a. The probability that an observation will fall between 240 and 300.
z(240) = (240-240)/40 = 0
z(300) = (300-240)/40 = 3/2
P(240< x <300) = P(0< z < 3/2) = 0.4332
-----
Fallow the same procedure for b,c, and d.
Cheers,
Stan H.
b. The probability that an observation will fall between 195 and 240.
c. The probability that an observation will be less than 260.
d. The probability that an observation will fall between 180 and 220.